# A Comprehensive Analysis of Social Tie Strength: Definitions, Prediction Methods, and Future Directions

### v2

@@
Abstract
The rapid growth of online social networks has underscored the importance of understanding the intensity of user relationships, referred to as “tie strength.” Over the past few decades, extensive efforts have been made to assess tie strength in networks. However, the lack of ground-truth tie strength labels and the differing perspectives on tie strength among researchers have complicated the development of effective prediction methods for real-world applications. In our study, we first categorize mainstream understandings of tie strength into seven standardized definitions and verify their effectiveness by investigating the class distributions and correlations across these definitions. We also draw key insights into tie resilience from the perspective of tie dissolution that (1) stronger ties are more resilient than weaker ones, and (2) this tie resiliency ratio increases as the network evolves. We then conduct extensive experiments to evaluate existing tie strength prediction methods under these definitions, revealing that (1) neural network methods capable of learning from semantic features hold great potential for high performance, (2) models struggle under definitions that offer limited understandings of tie strength in the network, (3) existing models face imbalance issues that cannot be addressed by traditional quantity imbalance techniques, and (4) different definitions of tie strength allow for the inference of not only the current state but also the future state of a tie. Building on these findings, we propose strategies to improve existing methods and suggest several promising directions for future research.
1 Introduction
@@
Recent years have witnessed the explosive growth of social networks. One of the most important components of social networks is the relationship between users, and the intensity of these relationships is often referred to as “tie strength” (Marsden and Campbell 1984). Tie strength has been widely studied and can have various definitions depending on the context and interpretation. For example, the tie strength between two users on Twitter (X) may be considered high if there is a bidirectional relation between them as it would indicate mutual interest. In another definition, two users may have a high tie strength if they frequently “like” or interact with each other’s posts on the platform. Extensive research has demonstrated the importance of understanding tie strength within social networks as it can extract meaningful insights from network data (Sintos and Tsaparas 2014). For instance, tie strength is a crucial factor in link prediction within networks (Li et al. 2013; Yang et al. 2019). Stronger ties in business networks are often correlated with successful trades, leading to better financial understanding (Kang et al. 2021; Wong et al. 2022). Additionally, tie strength helps predict epidemic and pandemic spread, aiding in disease prevention and improving healthcare planning (Krendl and Perry 2021; Birditt et al. 2021).
@@
During the past few decades, methods to infer/predict tie strength in social networks have been studied. The initial research (Granovetter 1973; Marsden and Campbell 1984; Petróczi, Nepusz, and Bazsó 2007; Gilbert and Karahalios 2009) developed heuristic methods to measure tie strength based on feedback from interviewees. Later researchers began to investigate using non-structural indicators such as frequency and duration of interaction (Zhuang et al. 2021; Jones et al. 2013; Xiang, Neville, and Rogati 2010), level of interaction or emotional intensity and intimacy (Gilbert and Karahalios 2009; Liu, Li, and Wang 2020), and the reciprocal services found within the tie (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022). However, these approaches generally rely on social network user characteristics that are often incomplete and difficult to measure (Adriaens et al. 2018). To address this issue, several methods have been developed that either rely solely on the Strong Triadic Closure (STC) principle (Sintos and Tsaparas 2014; Adriaens et al. 2018; Oettershagen, Konstantinidis, and Italiano 2022) or use heuristic approaches based on topological properties such as node degrees (Kahanda and Neville 2009), clustering coefficients (Mattie et al. 2018), or the overlap between nodes’ neighborhoods (Onnela et al. 2007) to characterize tie strength. In recent years, with the emergence of network embedding techniques, methods that generate embeddings to reflect the latent network structures properties have also been proposed (Zhong et al. 2020; Zhuang et al. 2021).
@@
However, ground-truth tie strength label information is usually missing or incomplete in practice. Most online social network platforms lack mechanisms to distinguish between different types of ties. Furthermore, obtaining ground-truth labels of tie strengths in networks is an under-explored problem (Liu, Li, and Wang 2020). Existing literature has various tie strength definitions, including direction-based definitions (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022), score-based definitions (Liu, Li, and Wang 2020) and frequency-based definitions (Zhuang et al. 2021). Since the differences between these definitions vary greatly, the tie strength prediction performance of these methods under different definitions also differs significantly. Therefore, it is necessary to devise a uniform definition of ground-truth tie strength labels and develop corresponding methods. As the field evolves, consistency across studies will enable more meaningful comparisons and allow future research to build on existing work more effectively.
As suggested by the empirical results later Section 5, existing models tend to have low prediction performance and insignificant average $tie$ weight differences between strong and weak ties. This can significantly affect the models’ ability to generalize across various real-world scenarios, and limit their effectiveness in diverse settings. In addition, failure to correctly recognize strong and weak ties can have significant implications in real-world networks, such as underestimating the influence of fraudulent behavior, or miscalculating the risk of a community suffering from a disease. It is important to understand the reasons why current tie strength prediction models fail to perform well under various tie strength definitions, and consider how we can obtain better predictions under weakly-/un-supervised settings.
To tackle these challenges, we first summarize the mainstream understandings of tie strength into seven standardized definitions and validate their significance by examining class distributions and correlations. Additionally, we draw valuable insights into tie resilience from the perspective of tie dissolution: (1) strong ties are more resilient than weak ties, and (2) this tie resiliency ratio increases as the network evolves. We then carried out extensive experiments to evaluate the performance of existing methods, leading to several key findings: (1) neural network models that effectively use semantic features have strong potential for high performance, (2) models face difficulties when using definitions that offer limited insights into tie strength, (3) current methods struggle with imbalance issues that cannot be addressed by traditional techniques, and (4) different definitions of tie strength allow for the prediction of different states of that tie in a dynamic network. Based on these findings, we propose strategies to improve existing models and suggest several promising directions for future work in tie strength prediction. The contributions of this study are as follows:
  • We summarize the mainstream practices for assigning tie strength labels and develop 7 standardized definitions, which we use as pseudo ground-truth tie strength labels.
  • We analyze the tie strength definitions, including correlation analysis and resilience of strong/weak ties from the perspective of tie dissolution.
  • We conduct experiments on tie strength prediction methods across numerous settings and develop new understandings from various perspectives.
  • We propose improvement strategies based on our experiments to highlight multiple future research directions.
The paper is structured as follows: We review existing tie strength prediction works and formalize the problem settings in Section 2. We revisit mainstream understandings of tie strength, and propose and investigate seven standardized definitions, including from a tie resilience perspective in Section 3. In Section 4, we present real-world settings and prediction methods. We provide extensive experiments, analyzing model performance across definitions and proposing strategies for improvement in Section 5. Finally, Section 7 concludes the paper.
2 Preliminaries
@@
2.1 Related Work
With the development of online social networks, tie strength prediction has received increasing attention  (Marsden and Campbell 2012; Gilbert 2012). While recent papers (Adriaens et al. 2018; Mattie et al. 2018) have explored tie strength at fine-grained levels, the most fundamental and widely accepted understanding (Granovetter 1973; Gilbert and Karahalios 2009; Kumar, Rao, and Nagpal 2012) favors the binary categorization of ties into strong/weak categories.
However, there is no standard way to obtain the ground truth tie strength (Liu, Li, and Wang 2020), and most online platforms even lack mechanisms to distinguish between different types of ties. To fill this gap, some of the pioneering work (Petróczi, Nepusz, and Bazsó 2007) relies on the feedback from questionnaires based on the theory proposed by  (Granovetter 1973), that is tie strength is a combination of the amount of time, the emotional intensity, the intimacy (mutual confiding) and reciprocal services.
Some other works also use semantic features to infer tie strength. For example,  (Xiang, Neville, and Rogati 2010) proposed a model to infer relationship strength based on profile similarity and interaction activity. However, such approaches rely heavily on the feedback qualities of users and, hence are hard to use for large social networks. To solve the issues, recent researchers began to develop heuristic methods based on structural features. For example,  (Mattie et al. 2018) developed heuristic methods that used structure features such as degree, clustering coefficients, unweighted overlap, etc as predictors.  (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022) also use strong triad closure properties (Granovetter 1973) to infer tie strength. Additionally, with the development of neural networks,  (Liu, Li, and Wang 2020; Zhuang et al. 2021) developed network embedding techniques to infer tie strength.
Despite these efforts, existing research poses different perspectives towards strong/weak ties. Some researchers  (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022) built their methods based on directional-based definitions that assume bidirectional edges are strong ties, while other researchers (Zhuang et al. 2021; Liu, Li, and Wang 2020; Zhuang et al. 2021; Jones et al. 2013; Xiang, Neville, and Rogati 2010) use score or frequency criteria to determine the ground-truth tie strength. To the best of our knowledge, no previous research has investigated these understandings and how different models perform under various perspectives, which makes it challenging to ensure solid performance from existing models in different real-world scenarios. To deepen our understanding of tie strength and shed light on future directions, we explicitly conduct this paper to comprehensively study the performance of existing tie strength methods under different tie strength definitions.
2.2 Notations
@@
The diagonal matrix of node degree is notated as $\mathbf{D}\in\mathbb{Z}^{n\times n}$ with the degree of node $v_{i}$ being $d_{i}=\mathbf{D}_{ii}=\sum_{j=1}^{n}\mathbf{A}_{ij}$. The neighborhood node set of node $v_{i}$ is denoted as $\mathcal{N}_{i}$.
2.3 Tie Strength Prediction Problem Settings
@@
In the tie strength prediction tasks, let the (pseudo) ground-truth tie strength label be $\mathbf{Y}\in\{0,1\}^{n\times n\times C}$, where $C$ is the total number of discrete tie strength categories, and $\{\mathcal{E}_{k}\}_{k=1}^{C}$ is the set of labeled ties with category $k$. Note that since tie strength is a reflection of the intensity of reciprocal relationships between users, a tie is undirected, which means there is only one tie strength between node $v_{i}$ and node $v_{j}$. Then:
@@
  • Weakly-supervised settings: Given a directed network $G$ with training tie strength set $\mathbf{Y}^{\text{Tr}}$, the task aims to correctly infer the strength of ties between nodes in the $\mathbf{Y}^{\text{Test}}$.
@@
  • Unsupervised settings: Given a directed network $G=$ with no training tie strength set, the task aims to correctly infer the strength of ties between nodes in the $\mathbf{Y}$.
Note that in this study we only consider binary tie strength (i.e., $C=2$) where existing edges can be either strong ($S$) or weak ($W$), which is a common setting in this line of research (Liu, Li, and Wang 2020).
3 Investigating Tie Strength Definitions
@@
In this section, we start by analyzing the existing tie strength definitions and organize them into 3 main categories. Then we propose a set of 7 tie strength definitions that span the primary tie strength categories from existing literature. To develop a more comprehensive understanding of these definitions, we collect three real-world datasets to conduct various analyses. Specifically, this includes analyzing the strong/weak tie distributions, pairwise correlation of strong/weak labeling of edges between definitions, and the resilience of ties from the perspective of tie dissolution.
3.1 Tie Strength Definitions
@@
The early work related to tie strength often relies on feedback from users who agree to participate in a survey (Petróczi, Nepusz, and Bazsó 2007; Gilbert and Karahalios 2009). However, this approach heavily relies on the feedback quality and is hard to generalize to large-scale social networks. In recent years, researchers have started to develop different methods to assign labels of tie strength based on different application scenarios. Most of the efforts in this direction can be roughly categorized into three categories:
@@
  • Direction-based Definitions: As suggested in  (Sintos and Tsaparas 2014; Adriaens et al. 2018; Zhong et al. 2020; Oettershagen, Konstantinidis, and Italiano 2022), a “strong tie” can be understood as a bidirectional/reciprocal relationship between users, as it would imply that the users know each other.
@@
  • Score-based Definitions: Another approach is to assign tie strength labels based on the score between users. For example,  (Liu, Li, and Wang 2020) proposed assigning tie strength based on rating scores (which represent the level of trust between buyers/sellers) exchanged between anonymous users in the Bitcoin networks.
@@
  • Frequency-based Definitions: Many works also assign tie strength labels based on frequency-related information. For instance,  (Zhuang et al. 2021; Jones et al. 2013; Xiang, Neville, and Rogati 2010) all consider tie strength through the lens of how frequently users interact.
@@
In most real-world datasets, both score-related and frequency-related information can usually be reflected in edge weight (i.e., $\mathbf{W}$). For example, the edge weight in Bitcoin_alpha (Derr et al. 2017) datasets represents the trust level rating score from one user to another, and the edge weight in the CollegeMSG (Panzarasa, Opsahl, and Carley 2009) dataset represents the contact frequency from one user to another. Hence, we summarized the existing understandings of tie strength into 7 definitions based on the edge weight and edge direction between users. More specifically, strong edges (i.e., $\mathbf{Y}_{ijS}=1$) are assigned according to the following conditions:
@@
Table 1: Comparison among definitions of strong ties based on their bidirectionality and weight information.
Tie Strength Definitions

### v5

# BTS: A Comprehensive Benchmark for Tie Strength Prediction

@@
Abstract.
The rapid rise of online social networks underscores the need to understand the heterogeneous strengths of online relationships. Yet, efforts to assess tie strength (TS) are hindered by the lack of ground-truth labels, differing research perspectives, and limited model performance in real-world settings. To address this gap, we introduce BTS, a comprehensive Benchmark for Tie Strength prediction, aiming to establish a standardized foundation for evaluating and advancing TS prediction methodologies. Specifically, our contributions are: TS Pseudo-Label Techniques—we categorize TS into seven standardized pseudo-labeling techniques based on prior literature; TS Dataset Collection—we present a representative collection of three social networks and perform data analysis by investigating the class distributions and correlations across the generated pseudo-labels; TS Pseudo-Label Evaluation Framework—we propose a standardized framework to evaluate the pseudo-label quality from the perspective of tie resilience; Benchmarking—we evaluate existing tie strength prediction model performance using the BTS dataset collection, exploring the effects of different experiment settings, models, and evaluation criteria on the results. Furthermore, we derive key insights to enhance existing methods and shed light on promising directions for future research in this domain. The BTS dataset collection, along with the curation codes and experimental scripts, is all available at: https://github.com/XueqiC/Awesome-Tie-Strength-Prediction.
Tie strength prediction; social network analysis; social tie resilience
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737441isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Social networks
1. Introduction
@@
Social networks have grown rapidly in recent years, becoming central to online interaction. A key component of these networks is the relationship between users, where the intensity of a connection—commonly referred to as tie strength (TS) (Marsden and Campbell, 1984)—has been widely studied. However, its definition can vary by context, application, or interpretation, and does not necessarily have a unified definition; for instance, a tie may be considered strong if it is bidirectional, indicating mutual interest, or may be inferred when users frequently engage with each other’s content. Understanding TS is critical, as it enables meaningful insights from social network data (Sintos and Tsaparas, 2014). For example, knowledge of TS can improve link prediction (Li et al., 2013; Yang et al., 2019). In business networks, strong ties often correlate with successful trades and better financial outcomes (Kang et al., 2021; Wong et al., 2022), while in public health, TS supports modeling disease spread, aiding prevention strategies, and healthcare planning (Krendl and Perry, 2021; Birditt et al., 2021).
@@
During the past decades, social tie strength prediction/inference methods have been studied. The initial research developed heuristic methods to measure tie strength based on feedback from interviewees (Granovetter, 1973; Marsden and Campbell, 1984; Petróczi et al., 2007; Gilbert and Karahalios, 2009). Later, researchers began to investigate using non-structural indicators such as frequency and duration of interaction (Zhuang et al., 2021; Jones et al., 2013; Xiang et al., 2010), level of interaction or emotional intensity and intimacy (Gilbert and Karahalios, 2009; Liu et al., 2020), and the reciprocal services found within the tie (Sintos and Tsaparas, 2014; Adriaens et al., 2018; Zhong et al., 2020; Oettershagen et al., 2022). However, these approaches generally rely on social network user characteristics that are often incomplete and difficult to measure (Adriaens et al., 2018). To address this issue, several methods have been developed that either rely solely on the Strong Triadic Closure (STC) principle (Sintos and Tsaparas, 2014; Adriaens et al., 2018; Oettershagen et al., 2022) or use heuristic approaches based on topological properties such as node degrees (Kahanda and Neville, 2009), clustering coefficients (Mattie et al., 2018), or the overlap between nodes’ neighborhoods (Onnela et al., 2007) to characterize tie strength. In recent years, methods for generating embeddings that capture latent network structures have also emerged and gained attention (Zhong et al., 2020; Zhuang et al., 2021).
@@
However, ground-truth tie strength label information is usually missing or incomplete in practice. Most online social network platforms lack mechanisms to distinguish between different types of ties. Furthermore, obtaining ground-truth labels of tie strengths in networks is an under-explored problem (Liu et al., 2020). Existing literature has various tie strength definitions, including direction-based definitions (Sintos and Tsaparas, 2014; Adriaens et al., 2018; Zhong et al., 2020; Oettershagen et al., 2022), score-based definitions (Liu et al., 2020), and frequency-based definitions (Zhuang et al., 2021). Since the differences between these definitions vary greatly, the empirical tie strength prediction performance of these methods under different definitions also differs significantly. We therefore need standardized pseudo-label techniques to summarize these definitions to ensure consistency across studies, enabling more meaningful comparisons and allowing future research to build upon existing work more effectively.
Furthermore, as suggested by the empirical results in Section 7, existing models exhibit low predictive performance and inconsistent accuracy between strong and weak ties, which can severely impact their ability to generalize across diverse real-world scenarios, limiting their effectiveness in applications such as detecting fraudulent behavior or assessing the risk of disease outbreaks within communities. Therefore, it is crucial to establish a benchmark to comprehensively evaluate and compare the performance of current and future tie strength prediction models across different settings.
To tackle these challenges, we present BTS, a comprehensive Benchmark for Tie Strength prediction, aiming to establish a standardized foundation for evaluating and advancing tie strength prediction methods. Our main contributions are as follows:
  • TS Pseudo-Label Techniques: We categorize mainstream understandings of social tie strength into seven standardized pseudo-label definitions that assign each edge as strong or weak;
  • TS Dataset Collection: We present a well-curated collection of three social networks, offering concise representations of diverse networks that capture different real-world social interactions. We also perform dataset analysis by investigating the class distributions and correlations across the generated pseudo-labels;
  • TS Pseudo-Label Evaluation Framework: A standardized framework is developed to evaluate TS pseudo-labeling, assuming strong ties are more resilient than weak ties over time.
  • Benchmarking: We evaluate existing tie strength prediction model performance using the BTS dataset collection, exploring the effects of different experiment settings, models, and evaluation criteria on the results. Furthermore, we derive key insights to enhance existing methods and shed light on promising directions for future research in this domain.
The paper is structured as follows: We review existing tie strength prediction works and formalize the problem settings in Section 2. We revisit mainstream understandings of tie strength and propose standardized pseudo-label techniques in Section 3. In section 4, we introduce the TS dataset collection and perform empirical analysis based on it using the proposed pseudo-label techniques. In Section 5, we further develop the tie resilience evaluation framework. In Section 6, we present a comprehensive benchmark of existing tie strength prediction methods. Section 8 discusses the limitations and future work. Finally, Section 9 concludes the paper.
2. Preliminaries
@@
2.1. Related Work
With the development of online social networks, tie strength prediction has received increasing attention  (Marsden and Campbell, 2012; Gilbert, 2012). While recent papers (Adriaens et al., 2018; Mattie et al., 2018) have explored tie strength at fine-grained levels, the most fundamental and widely accepted understanding (Granovetter, 1973; Gilbert and Karahalios, 2009; Kumar et al., 2012) favors the binary categorization of ties into strong/weak categories.
However, there is no standard way to obtain the ground truth tie strength (Liu et al., 2020), and most online platforms even lack mechanisms to distinguish between different types of ties. To fill this gap, some of the pioneering work (Petróczi et al., 2007) relies on the feedback from questionnaires based on the theory proposed by  (Granovetter, 1973), that is, tie strength is a combination of the amount of time, the emotional intensity, the intimacy (mutual confiding) and reciprocal services.
Some other works also use semantic features to infer tie strength. For example,  (Xiang et al., 2010) proposed a model to infer relationship strength based on profile similarity and interaction activity. However, such approaches rely heavily on the feedback qualities of users and are hard to use for large social networks. To solve the issues, recent researchers have begun to develop heuristic methods based on structural features. For example,  (Mattie et al., 2018) developed heuristic methods that used structure features such as degree, clustering coefficients, unweighted overlap, etc, as predictors.  (Sintos and Tsaparas, 2014; Adriaens et al., 2018; Zhong et al., 2020; Oettershagen et al., 2022) also use strong triad closure properties (Granovetter, 1973) to infer tie strength. Additionally, with the development of neural networks,  (Liu et al., 2020; Zhuang et al., 2021) developed network embedding techniques to infer tie strength.
Despite these efforts, existing research poses different perspectives towards strong/weak ties. Some researchers (Sintos and Tsaparas, 2014; Adriaens et al., 2018; Zhong et al., 2020; Oettershagen et al., 2022) built their methods based on directional-based definitions that assume bidirectional edges are strong ties, while other researchers (Zhuang et al., 2021; Liu et al., 2020; Zhuang et al., 2021; Jones et al., 2013; Xiang et al., 2010) use score or frequency criteria to determine the ground-truth tie strength. To the best of our knowledge, no previous research has investigated these understandings and how different models perform under various perspectives, which makes it challenging to ensure solid performance from existing models in different real-world scenarios. To deepen our understanding of tie strength and shed light on future directions, we explicitly conduct this paper to comprehensively study the performance of existing tie strength methods under different tie strength definitions.
2.2. Notations
@@
2.3. Tie Strength Prediction Problem Settings
@@
In the tie strength prediction tasks, let the (pseudo) ground-truth tie strength label be $\mathbf{Y}\in\{0,1\}^{n\times n\times C}$, where $C$ is the total number of discrete tie strength categories, and $\{\mathcal{E}_{k}\}_{k=1}^{C}$ is the set of labeled ties with category $k$. Given that tie strength is a reflection of the intensity of reciprocal relationships between users, a tie is undirected, which means there is only one tie strength between node $v_{i}$ and node $v_{j}$. Note that in this study, we only consider binary tie strength (i.e., $C=2$) where existing edges can be either strong ($S$) or weak ($W$), which is a common setting in this line of research (Liu et al., 2020). Then, the weakly-supervised and unsupervised TS prediction settings can be defined as follows:
@@
  • Weakly-supervised settings: Given a directed network $G$ with training tie strength set $\mathbf{Y}^{\text{Tr}}$, the task aims to correctly infer the strength of ties between nodes in $\mathbf{Y}^{\text{Test}}$.
@@
  • Unsupervised settings: Given a directed network $G$ with no training tie strength set, the task aims to correctly infer the strength of ties between nodes in $\mathbf{Y}$.
3. Tie Strength Pseudo-Label Techniques
@@
In this section, we start by analyzing the existing tie strength definitions and organizing them into 3 main categories. Then, we propose a set of 7 tie strength definitions that span the primary tie strength categories from existing literature. To develop a more comprehensive understanding of these definitions, we collect three real-world datasets to conduct various analyses. Specifically, this includes analyzing the strong/weak tie distributions, pairwise correlation of strong/weak labeling of edges between definitions, and the resilience of ties from the perspective of tie dissolution.
@@
The early work related to tie strength often relies on feedback from users who agree to participate in a survey (Petróczi et al., 2007; Gilbert and Karahalios, 2009). However, this approach heavily depends on the quality of feedback and is hard to generalize to large-scale social networks. In recent years, researchers have developed different methods to assign labels of tie strength based on different application scenarios. Most of the efforts in this direction can be roughly categorized into three categories:
@@
  • Direction-based Definitions: As suggested in  (Sintos and Tsaparas, 2014; Adriaens et al., 2018; Zhong et al., 2020; Oettershagen et al., 2022), a “strong tie” can be understood as a bidirectional/reciprocal relationship between users, as it implies users know each other.
@@
  • Score-based Definitions: Another approach is to assign tie strength labels based on the score between users. For example,  (Liu et al., 2020) proposed assigning tie strength based on rating scores (which represent the level of trust between buyers/sellers) exchanged between anonymous users in the Bitcoin networks.
@@
  • Frequency-based Definitions: Many works also assign tie strength labels based on frequency-related information. For instance,  (Zhuang et al., 2021; Jones et al., 2013; Xiang et al., 2010) all consider tie strength through the lens of how frequently users interact.
@@
In most real-world datasets, both score-related and frequency-related information can usually be reflected in edge weight (i.e., $\mathbf{W}$). For example, the edge weight in Bitcoin_alpha (Derr et al., 2017) datasets represents the trust level rating score from one user to another, and the edge weight in the CollegeMSG (Panzarasa et al., 2009) dataset represents the contact frequency from one user to another. Hence, we summarized the existing understandings of tie strength into standardized pseudo-labels based on the edge weight and edge direction between users. More specifically, strong edges (i.e., $\mathbf{Y}_{ijS}=1$) are assigned according to the following conditions:
@@
Table 1. Comparison among different pseudo-labels of strong ties based on their bidirectionality and weight information.
Edge Conditions Pseudo-labeling Techniques